1)Which one of the fowing statements is not cnereet b) Displacement is a vector quintity. c) Kinetic energy is a scalar quantity d) Coeficient of static Siction is a scalar quantity e) Impulse is a vector quantity 2) The unit of momennm is a) Nkgn b) kgms c)Nkgs e) None of the ahove 8) The x coordinate of a particle moving along the x-axis at different times is given in the table. Wha ¡s the particles avenge velocity in ms between t-2sandt 5s? 12 2 36 43 60 28492 a) 92 b) 72 c) 48 d) 36 e) 12 4) A ball is thrown with an initial velocity of ve- 10 ms at angle 8-53, from the top of a buildit of height h-30 m. The vertical displacement (in m) of the ball when it hits the ground is: a) 30 b) 0 c)-30 d)-24 e) -100 田田| 5) A stone is thrown vertically upward. If we neglect air resistance, as it rises, its total mechan energy (kinetic energy + potential energy): a) increases b) decreases c) is constant d) decreases then increases e) increases then decreases

please ask the professors to answer the questions quickly because this assignment must be delivered within one and a half hours
Answer #1

1) Momentum is a vector quantity, therefore statement (a) is incorrect.

2) The unit of momentum is kgm/s and N.s. Therefore (b) and (d) are correct.

3) The average velocity = x5-x2/(5-2) = 60-24/3 = 12 m/s.

4) The vertical displacement is - h = -30 m. Option c. The net displacement above the building is zero (H-H =0).

5) The kinetic energy + potential energy (mechanical energy) is constant as air resistance is 0. Option c is correct.